North: Pat O. had a team of Pat St, Lori and Michael a storm-threatening morning. They had a U turn false crawl and a nest that needed relocation.
DEB29 Relocated north #10mm laid below spring tide line Lori located!
DEBFC31 north #14mm
Hobcaw: Duane delayed his walk until around 11:00 when the weather improved. He had a total of 4 crawls before he turned around at #16mm due to more weather coming. Gimpy came back yesterday (HOBFC45) which we didn’t recognize as her crawl. She last nested 6/14 so she is right on time and successfully nested today:) Survey well done today, Duane!
HOB22 Natural south #9mm (Gimpy)
HOB23 Natural south #5mm in grasses (barnacle on plastron)
HOBFC48 south #7mm (Gimpy)
HOBFC49 north #6mm (Gimpy)
DEB 29 HOB 23 SCUTE 140 SC 4819
